OpenStudy (anonymous):

A card is selected from a standard deck of 52 playing cards. A standard deck of cards has 12 face cards and four Aces (Aces are not face cards). Find the probability of selecting · an odd prime number under 10 given the card is a club. (1 is not prime.) · a Jack, given that the card is not a heart. · a King given the card is not a face card. Show step by step work. Give all solutions exactly in reduced fraction form.

OpenStudy (anonymous):

given the card is a club means you have 13 in your sample space. the odd primes are 3, 5, 7

OpenStudy (anonymous):

so answer to first one is \[\frac{3}{13}\]

OpenStudy (anonymous):

not a heart is 39 cards and there are 3 jacks that are not hearts so it is \[\frac{3}{39}=\frac{1}{13}\]

OpenStudy (anonymous):

both of those are independent so the "given" part is irrelevant

OpenStudy (anonymous):

a king given that the card is not a face card is a joke. not possible so that answer is 0

OpenStudy (anonymous):

Ok Thanks